OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

solve the brackets first

OpenStudy (anonymous):

in the case of \[2(x+6)\]you just multiply 2 with x and 6 so it becomes 2x+12

OpenStudy (anonymous):

same applies with 4(x-1), this becomes 4x-4

OpenStudy (anonymous):

now we just group like terms 2x+12-10=12+4x-4 12-10-12+4=4x-2x

OpenStudy (anonymous):

-6=2x x=-6/2

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yes ur a great help

OpenStudy (anonymous):

I do what I can